OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x2 + 24x + 16? Answer (2, 8) (2, –8) (–2, –8) (–2, 8)

jhonyy9 (jhonyy9):

y=6(-2)^2 +24(-2) +16 =24-48+16 =40-48 =-8 so than (-2,-8) will be right